OpenStudy (anonymous):

Each of these sequence or series formulas involves four quantities, for example, a1, an, d, r, n, and Sn. For each formula, describe the four quantities. Then explain how you can find the fourth quantity if you know the values of the other three. A) an=a1 + (n-1) d B) sn= n/2 ( a1-an) C) an=a1r ^ n-1 D) Sn = a1 ( 1-r^n) / 1-r Dont know what to do..anyone know the answer?

OpenStudy (anonymous):

\(a_1\) stands for the first number in the sequence/ series. For example: in sequence 1,2,3,4. \(a_1 \) =1 1/\(a_n\) stands for the \(n^{th}\) number in the sequence/ series. For example : in the sequence above \(a_2 \) is \(2^{nd}\) number and \(a_2 =2\) 2/d stands for the difference between 2 consecutive numbers in the sequence. For example: int the sequence above \(a_2 =2, a_1=1\rightarrow d = a_2-a_1 =2-1=1\) 3/r stands for the ratio if the sequence is geometric one. For example: in sequence 2,4,8, 16... the ratio r = \(\dfrac{next~term}{previous~term}=\dfrac{a_2}{a_1}=\dfrac{4}{2}=2\) 4/n stands for numbers of the sequence. For example:2,4,8,16. you have n =4( because you have 4 elements in the sequence) 5/\(S_n\) stands for sum of the element in the sequence. For example: in sequence 1,2,3,4,5. \(S_5= a_1+a_2+a_3+a_4+a_5=1+2+3+4+5= 15\) They can ask you to find \(S_3\). That means you have to find sum of first 3 elements in the sequence \(S_3= 1+2+3=6\) Read and tell me whether you got the information or not. You Must understand the concept before solving your problem.